OpenStudy (anonymous):

The function H described by H(x) = 2.75x = 71.48 can be used to predict the height, in centimeters, of a woman whose humerus (the bone from teh elbow to the shouldber) is x cm long. Predict the height of a woman whose humerus is 33 cm long. The predicted height of a woman whose humerus is 33 cm long is ? cm I have 174.48 as the answer?

OpenStudy (anonymous):

I mean 160.48

OpenStudy (pfenn1):

The function is\[ H(x) =2.75x\]Is this the right equation or is something missing?

OpenStudy (anonymous):

I have H(33)=2.75+71.48

OpenStudy (pfenn1):

I was wondering.....that isn't what it showed above. so isn't the general equation given by\[H(x) =2.75x+71.48\]?

OpenStudy (anonymous):

When I reentered the problem I got 162.23?

OpenStudy (pfenn1):

Well if the equation is as I last wrote it then\[H(33) =2.75(33)+71.48=162.23\]I agree.